OpenStudy (anonymous):

The whiteboard in the SPA has a perimeter of 408 inches. The length is 11 inches less than 4 times the height. How many inches high is the whiteboard?

OpenStudy (mathstudent55):

Let L = length and H = height The perimeter of a rectangle is the sum of the lengths of all the sides. P = L + L + H + H 408 = 2L + 2H 2L + 2H = 408 (this is one equation) The following statement will gives us the other equation. The length is 11 inches less than 4 times the height. L = 4H - 11 L = 4H - L (this is the second equation) Now you need to solve the system of equations: 2L + 2H = 408 L = 4H - 11